# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ II.1 Dynamical constraints and semigroups
where $G_{\ell_{R}}$ denotes the set of elements of $G$ expressible as products of precisely $\ell_{R}$ generators. A particularly natural realization of ${\sf Dyn}_{G}$ is when each $U_{g}$ is drawn from an appropriate-dimensional Haar ensemble. Existing examples of constrained 1D dynamics in the literature—from multipole conserving systems to models based on cellular automata and other constrained classical systems Sala et al. (2020b); Khemani et al. (2020b); Moudgalya and Motrunich (2022); Pancotti et al. (2020); van Horssen et al. (2015); Brighi et al. (2023); Lan et al. (2018); Garrahan (2018)—are all described by ${\sf Dyn}_{G}$ for an appropriate semigroup $G$ and an appropriate kind of dynamics (random unitary, Hamiltonian, etc.). [^5] This is not true for models with quantum Hilbert space fragmentation Moudgalya and Motrunich (2022); Brighi et al. (2023), which violate the above assumptions by virtue of having constraints that cannot be formulated in a local product basis. One may then hope to develop a general theory of the different dynamical universality classes that can arise in constrained 1D systems by using results on the structure and classification of semigroups. This ambition is damped somewhat by the fact that semigroups are very general mathematical objects, and this generality limits our ability to make concrete statements. One of the main points of this work however is that significantly more progress can be made when $G$ is a group, viz. a semigroup with inverses and an identity. Making this restriction allows a large arsenal of mathematical tools from the field of geometric group theory to be brought to bear, leading to general bounds on thermalization times, precise characterizations of ergodicity breaking, and so on.
@@ II.2 The word problem
We will now formulate the semigroup word problem, a concept key for determining the thermalization beahvior of models with semigroup dynamics. We will refer to a computational basis product state—defined by a string of generators in $S$, e.g. $w=\mathtt{a}_{1}\cdots\mathtt{a}_{L}$—as a word. Words are naturally grouped into equivalence classes labeled by elements of $G$. Letting $W(S)$ denote the set of all words, we define these equivalence classes as
@@
where each arrow $\to$ indicates applying a single relation from $R$. In Sec. III we will see that the way in which ${\sf Dyn}_{G}$ thermalizes is determined by the complexity of the word problem for $G$, a fundamental problem in the fields of abstract algebra and computability theory. The word problem is defined by the following question: Word problem: Given two words $w,w^{\prime},$ does $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$? That is, does there exist a derivation $D(w\rightsquigarrow w^{\prime})$? A key result is that even in the case where both $|S|$ and $|R|$ are small, answering this question can be very difficult (even undecidably so; see App. C). For semigroups or groups which do not have an undecidable word problem, a key problem is to determine the time and space complexity of algorithms that solve it. In what follows, we will introduce two functions characterizing the word problem’s complexity: the Dehn function, which governs its complexity, and the expansion length, which governs its space complexity.
@@ II.2.1 Time complexity
Our operational definition of the time complexity of the word problem is the minimum length of a derivation linking $w$ to $w^{\prime}$, as illustrated in Fig. 2 (a). [^6] An important caveat here is that proving $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$ does not need to require outputing an explicit derivation $D(w_{1}\rightsquigarrow w_{2})$, as for a particular $G$ there may exist other (faster) algorithms for constructing a proof. That is, when we speak of the time complexity of the word problem, we are referring to its time complexity within the so-called Dehn proof system, a logical system where the only manipulations one is allowed to perform are the implementation of local relations in $R$. If one wants an algorithm that is not tailored for a certain $G$, operating within this system is generically the best one can do. We denote this by ${\sf Dehn}(w_{1},w_{2})$:
@@
as the set of length-$L$ words in $K_{g}$. The (worst-case) time complexity across all words in $K_{g,L}$ defines the function
@@
We will be particularly interested in how ${\sf Dehn}_{g}(L)$ scales asymptotically with $L$. In App. B, we show that this scaling is the same for any two finite presentations of the same semigroup: thus we may meaningfully speak about the Dehn function of a semigroup, rather than a particular presentation thereof. This presentation independence implies some degree of the robustness of the dynamical properties that we will derive. For the case where $G$ is a group rather than just a semigroup, a fair amount more can be said. All groups have a distinguished identity element $e$, and in App. B we show that [^7] Throughout this work, ,$\sim$,$\gtrsim$,$\lesssim$ , ≳ , ≲ will be used to denote (in)equality in the scaling sense. For example, when we write $f(L)\sim g(L)$, we mean that there exists a constant $C>0$ such that $f(L/C)\leq g(L)\leq f(LC)$.
@@
${\sf Dehn}(L)=\max_{|w\rangle\in K_{e,L}}{\sf Dehn}(w,\mathtt{e}^{L})\qquad\text{($G$ a group)}$) (17)
@@
with ${\sf Dehn}(L)\sim L^{2}$ only when $G$ is infinite. Indeed, for all such groups, such as the $G=\mathbb{Z}^{2}$ example above, ${\sf Dehn}(L)$ is bounded from above by the time it takes to transport the conserved charges $n_{\mathtt{a}}\triangleq\sum_{i}(|\mathtt{a}\rangle\langle\mathtt{a}|_{i}-|\mathtt{a}^{-1}\rangle\langle\mathtt{a}^{-1}|_{i})$ across the system, which is $\sim L^{2}$. For our purposes, we will regard any $G$ with ${\sf Dehn}(L)\lesssim L^{2}$ as uninteresting, as for these groups ${\sf Dyn}_{G}$ thermalizes on timescales generically no slower than for conventional systems with conserved $U(1)$ charges. A simple example of a group with an “interesting” Dehn function is the discrete Heisenberg group $\mathsf{H}_{3}=\langle\mathtt{x},\mathtt{y},\mathtt{z}\,|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\mathtt{z},\,\mathtt{x}\mathtt{z}=\mathtt{z}\mathtt{x},\,\mathtt{y}\mathtt{z}=\mathtt{z}\mathtt{y}\rangle$, which has ${\sf Dehn}(L)\sim L^{3}$ Gersten et al. (2003). In Sec. IV we will see an example of a simple group where ${\sf Dehn}(L)\sim 2^{L}$ and in Sec. V one with ${\sf Dehn}(L)\sim 2^{2^{L}}$. Going beyond these examples, Refs. Brady and Bridson (2000); Sapir et al. (2002) remarkably show that for any constant $\alpha$, almost any function with growth $L^{2}\leq f(L)<L^{\alpha}$ is the Dehn function of some finitely presented group; this includes for example “unreasonable-looking” functions such as $L^{7\pi},L^{2}\log L,$ and $L^{e}\log(L)^{\log(L)}\log\log L$. In addition to the worst case complexity of the word problem, we will also have occasion to consider its average-case complexity (both for groups and semigroups), a quantity has recieved much less attention in the math literature (the only exception the authors are aware of is Young (2008)). To this end, we define the typical Dehn function $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu_{g}(L)$ as the number of steps needed to map a certain constant fraction of words in $K_{g,L}$ to one another:
@@ II.2.2 Space complexity
${\sf EL}(L)\triangleq{\sf EL}_{e}(L)\qquad\text{($G$ a group)}$) (23)
as a simple metric of the space complexity. In App. B we show that ${\sf EL}(L)\lesssim L$ for all Abelian groups, and that all finite groups have ${\sf EL}(L)\leq L+C$ for some constant $C$. Groups with ${\sf EL}(L)\lesssim L$ are thus “uninteresting” from the perspective of space complexity. Just as with the Dehn function though, there exist simple groups for which ${\sf EL}(L)$ grows extremely fast with $L$, the consequences of which will be explored in Sec. V.
@@ II.3 Semigroup dynamics and Hilbert space fragmentation
The existence of multiple $K_{g,L}$ sectors means that ${\sf Dyn}_{G}$ is not ergodic as long as $G$ is not a presentation of the trivial group. This lack of ergodicity is true for both semigroups and groups, but for simplicity we will focus our attention to the case of groups. In the case where $G$ is Abelian, $K_{g,L}$ can be associated with the symmetry sectors of a global symmetry. This is true simply because the sector that a given product state lives in can be determined by computing the expectation value of the operators $n_{\mathtt{g}}=\sum_{i}(|\mathtt{g}\rangle\langle\mathtt{g}|_{i}-|\mathtt{g}^{-1}\rangle\langle\mathtt{g}^{-1}|_{i}])$ for each generator $\mathtt{g}$. Thus the dynamics of Abelian groups is well understood given the plethora of work on thermalization in the presence of global symmetries. In the (generic) case where $G$ is non-Abelian, global symmetries may still be present, but there inevitably exist other non-local conserved quantities which distinguish different dynamical sectors. Indeed, in App. D we prove that the dynamical sectors of non-Abelian group dynamics can never be described by global symmetries alone. Since ${\sf Dyn}_{G}$ with non-Abelian $G$ will thus always have non-local conserved quantities, the lack of ergodicity due to these quantities leads to Hilbert space fragmentation (HSF) Khemani et al. (2020c); Sala et al. (2020c); Moudgalya et al. (2022c); Moudgalya and Motrunich (2022, 2023), a phenomenon whereby the dynamics of initial states becomes trapped in disconnected subspaces of $\mathcal{H}$—in our notation simply the $K_{g,L}$—whose existence cannot be attributed to the presence of global symmetries alone. The original works on HSF Khemani et al. (2020c); Sala et al. (2020c) focused mainly on fragmentation in spin systems with conserved dipole moments. While we will revisit this example later, for now we will find it more instructive to review another example which is more directly linked to groups: the pair-flip model Caha and Nagaj (2018). The pair-flip model is described by the following spin-$s$ Hamiltonian:
@@
For generic choices of $\lambda_{ab,i}$, the model is non-integrable, so any ergodicity breaking cannot be associated with integrability. Note that the product states $|w\rangle$ where $w=a_{1}\cdots a_{L}$ are annihilated by $H_{PF}$ if $a_{i}\neq a_{i+1}$ for all $i$. These product states alone provide $(2s+1)(2s)^{L-1}$ dynamically-disconnected dimension-1 sectors not attributable to any local symmetry, meaning that $H_{PF}$ exhibits HSF. $H_{PF}$ is in fact a special case of our construction applied to the group
@@
which matches the Hamiltonian in (24). Unfortunately the Dehn function of $\mathbb{Z}_{2}^{\ast(2s+1)}$ is easily seen to scale as ${\sf Dehn}(L)\sim L^{2}$, and thus does not provide a complexity scaling that is unusual enough to warrant further study. Having discussed the concept of HSF, let us further understand the structure and size of disconnected sectors under the dynamics ${\sf Dyn}_{G}$. The number of such disconnected sectors at least as large as the number of subspaces, each labelled by $K_{g,L}$ for some $g\in G$. Thus, the number of such subspaces—which we denote as $N_{K}(L)$—depends on the number of group elements expressible as words of length $\leq L$. In particular, we define the geodesic length $|g|$ of an element $g\in G$ as the length of the shortest word representing $g$:
@@
A word $w\sim g$ satisfying $|w|=|g|$ is called a geodesic word. Then
@@
Groups with more elements “close” to the identity thus have dynamics with a larger number of Hilbert space fragments. One may readily verify that in the pair flip model, $N_{K}(L)\sim(2s)^{L}$ grows exponentially as long as $s>1/2$. For many models with group constraints, including the pair-flip model, $N_{K}(L)$ exactly determines the number of Hilbert space fragments. However, for some groups, $N_{K}(L)$ is not the full story, with each $K_{g,L}$ further fragmenting into subspaces in a way controlled by the expansion length function defined in (22). Understanding this phenomenon (which we call fragile fragmentation) is the subject of Sec. V. In the study of HSF, a distinction is often made between “strong” and “weak” HSF. This distinction was originally discussed in the context of Hamiltonian dynamics Sala et al. (2020b), where it was defined by violations of strong and weak ETH, respectively. Since we are discussing things at a level where the nature of ${\sf Dyn}_{G}$ may or may not involve eigenstates, we will instead adopt a slightly different definition in terms of the size of the largest $K_{g,L}$ sector, which we denote as $K_{{\rm max},L}$ (in App. B we prove that for groups, the largest sector is in fact always the one associated with the identity, $K_{{\rm max},L}=K_{e,L}$). We will say that the dynamics is
@@
We will find it useful to subdivide the strongly fragmented case into additional classes according to how quickly $|K_{e,L}|/|\mathcal{H}|$ vanishes as $L\rightarrow\infty$. We will say that the fragmentation in the case of strong HSF is
@@
Note that the above definitions are made without reference to any global symmetry sectors. This means that if global symmetries happen to be present, the quantum numbers associated with them will constitute part of the elements $g$ labeling the different dynamical sectors, and the above definition of strong/weak HSF will simply single out the largest, regardless of its symmetry quantum number(s). When symmetries are present one could also quantify the degree of fragmentation by first fixing a quantum number, changing $K_{\rm max}(L)$ to be the largest sector having that quantum number, and replacing $|\mathcal{H}|$ by the dimension of the chosen symmetry sector. However, since generic ${\sf Dyn}_{G}$ dynamics needn’t have any global symmetries, and since the result of the above procedure can depend sensitively on the chosen quantum number Morningstar et al. (2020); Pozderac et al. (2023); Wang and Yang (2023), we will focus only on the above (simpler) definition, which maximizes over all symmetry sectors. Our models provide a way of addressing two questions raised in Ref. Sala et al. (2020b). The first question was whether or not 1D models exist with $10<|K_{{\rm max},L}|/|\mathcal{H}|<$ 1 in the thermodynamic limit (known examples with weak fragmentation all have $|K_{{\rm max},L}|/|\mathcal{H}|\rightarrow 1$ in the thermodynamic limit). Our construction answers this question in the affirmative, with examples provided by ${\sf Dyn}_{G}$ for any finite non-Abelian $G$ (e.g. $G=S_{3}$). The second question concerned the existence of models where $|K_{{\rm max},L}|/|\mathcal{H}|$ vanishes more slowly than exponentially as $L\rightarrow\infty$ after specifying to a fixed symmetry sector (in fact an affirmative answer to this question was already provided by the spin-1 Motzkin chain introduced in Bravyi et al. (2012), where $|K_{{\rm max},L}|/|\mathcal{H}|\sim L^{-3/2}$). [^10] This follows from the fact that the probability of a length-$L$ random walk on $\mathbb{Z}$ being a returning Motzkin walk scales as $\sim L^{-3/2}$. Our models provide (many) more examples of this phenomenon, as one need only let $G$ be a group with $L^{2}\lesssim{\sf Dehn}(L)\lesssim L^{\infty}$, a simple example being the discrete Heisenberg group $\mathsf{H}_{3}$. [^11] If ${\sf Dyn}_{G}$ possesses global symmetries and one modifies the above definitions to operate within a fixed symmetry sector, then various models are known where a specific symmetry sector with polynomially strong fragmentation exists Morningstar et al. (2020); Pozderac et al. (2023); Wang and Yang (2023). One may further ask whether there are strongly fragmented systems where $|K_{{\rm max},L}|/|\mathcal{H}|$ decays at a rate in between $1/{\rm poly}(L)$ and $\exp(-L)$. The answer to this is again affirmative, with one such example being the focus of Sec. IV.

### 11 Apr 2024 v2

@@ II.1 Dynamical constraints and semigroups
where $G_{\ell_{R}}$ denotes the set of elements of $G$ expressible as products of precisely $\ell_{R}$ generators. A particularly natural realization of ${\sf Dyn}_{G}$ is when each $U_{g}$ is drawn from an appropriate-dimensional Haar ensemble, although we shall not need to specify to this case. Existing examples of constrained 1D dynamics in the literature—from multipole conserving systems to models based on cellular automata and other constrained classical systems [21, 22, 23, 24, 25, 26, 27, 28]—are all described by ${\sf Dyn}_{G}$ for an appropriate semigroup $G$ and an appropriate kind of dynamics (random unitary, Hamiltonian, etc.). [^5] This is not true for models with quantum Hilbert space fragmentation [23, 26], which violate the above assumptions by virtue of having constraints that cannot be formulated in a local product basis. One of the main messages of this paper is that in addition to providing an organizing framework for discussing 1D constrained dynamics, approaching things from this point of view enables a large arsenal of mathematical tools from the field of geometric group theory to be brought to bear, leading to general bounds on thermalization times, precise characterizations of ergodicity breaking, and explicit examples of models with extremely slow dynamics.
@@ II.2 The word problem
We will now formulate the semigroup word problem, a concept key for determining the thermalization beahvior of models with semigroup dynamics. We will refer to a computational basis product state—defined by a string of generators in $S$, e.g. $w=\mathtt{a}_{1}\cdots\mathtt{a}_{L}$—as a word. In what follows we will often slightly abuse notation by letting the symbol $w$ stand for both an abstract string $\mathtt{a}_{1}\cdots\mathtt{a}_{L}$ of generators in $S$, as well as the associated computational-basis product state $|\mathtt{a}_{1}\cdots\mathtt{a}_{L}\rangle$. Words are naturally grouped into equivalence classes labeled by elements of $G$. Letting $W(S)$ denote the set of all words, we define these equivalence classes as
@@
where each arrow $\to$ indicates applying a single relation from $R$. In Sec. III we will see that the way in which ${\sf Dyn}_{G}$ thermalizes is determined by the complexity of the word problem for $G$, a fundamental problem in the fields of abstract algebra and computability theory. The word problem is defined by the following question: Word problem: Given two words $w,w^{\prime},$ does $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$? That is, does there exist a derivation $D(w\rightsquigarrow w^{\prime})$? A key result is that even in the case where both $|S|$ and $|R|$ are small, answering this question can be very difficult (even undecidably so; see App. C). For semigroups or groups which do not have an undecidable word problem, a key problem is to determine the time and space complexity of algorithms that solve it. In what follows, we will introduce two functions characterizing the word problem’s complexity: the Dehn function, which governs its time complexity, and the expansion length, which governs its space complexity.
@@ II.2.1 Time complexity
Our operational definition of the time complexity of the word problem is the minimum length of a derivation linking $w$ to $w^{\prime}$, as illustrated in Fig. 2 (a). [^6] An important caveat here is that proving $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$ does not need to require outputing an explicit derivation $D(w_{1}\rightsquigarrow w_{2})$, as for a particular $G$ there may exist other (faster) algorithms for constructing a proof. We therefore must stress that when we speak of the time complexity of the word problem, we are referring to its time complexity within the so-called Dehn proof system, a logical system where the only manipulations one is allowed to perform are the implementation of local relations in $R$. If one wants an algorithm which works for all choices of $G$, operating within this system is generically the best one can do. We denote this by ${\sf Dehn}(w_{1},w_{2})$:
@@
as the set of length-$L$ words in $K_{g}$ (or equivalently, following our practice of letting $w$ stand for both a word and a computational-basis product state, as the collection of product states associated to such words). The (worst-case) time complexity across all words in $K_{g,L}$ defines the function
@@
We will be particularly interested in how ${\sf Dehn}_{g}(L)$ scales asymptotically with $L$. In App. B, we show that this scaling is the same for any two finite presentations of the same semigroup: thus we may meaningfully speak about the Dehn function of a semigroup, rather than a particular presentation thereof. This presentation independence imparts some degree of the robustness to the dynamical properties that we will derive. For the case where $G$ is a group (rather than just a semigroup), a fair amount more can be said. All groups have a distinguished identity element $e$, and in App. B we show that [^7] Throughout this work, ,$\sim$,$\gtrsim$,$\lesssim$ , ≳ , ≲ will be used to denote (in)equality in the scaling sense. For example, when we write $f(L)\sim g(L)$, we mean that there exists a constant $C>0$ such that $f(L/C)\leq g(L)\leq f(LC)$.
@@
${\sf Dehn}(L)=\max_{|w\rangle\in K_{e,L}}{\sf Dehn}(w,\mathtt{e}^{L})\qquad\text{({\hbox{G}} a group)}$) (17)
@@
with ${\sf Dehn}(L)\sim L^{2}$ only when $G$ is infinite. Indeed, for all such groups, such as the $G=\mathbb{Z}^{2}$ example above, ${\sf Dehn}(L)$ is bounded from above by the time it takes to transport the conserved charges $n_{\mathtt{a}}\triangleq\sum_{i}(|\mathtt{a}\rangle\langle\mathtt{a}|_{i}-|\mathtt{a}^{-1}\rangle\langle\mathtt{a}^{-1}|_{i})$ across the system, which is $\sim L^{2}$. For our purposes, we will regard any $G$ with ${\sf Dehn}(L)\lesssim L^{2}$ as uninteresting, as for these groups ${\sf Dyn}_{G}$ thermalizes on timescales generically no slower than for conventional systems with conserved $U(1)$ charges. A simple example of a group with an “interesting” Dehn function is the discrete Heisenberg group $\mathsf{H}_{3}$, which has the group presentation
$\mathsf{H}_{3}=\langle\mathtt{x},\mathtt{y},\mathtt{z}\,|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\mathtt{z},\,\mathtt{x}\mathtt{z}=\mathtt{z}\mathtt{x},\,\mathtt{y}\mathtt{z}=\mathtt{z}\mathtt{y}\rangle,$ (19)
and possesses a Dehn function scaling as ${\sf Dehn}(L)\sim L^{3}$ [42]. In Sec. IV we will see an example of a simple group where ${\sf Dehn}(L)\sim 2^{L}$ and in Sec. V one with ${\sf Dehn}(L)\sim 2^{2^{L}}$; examples of semigroups with similarly slow dynamics are given in Sec. VI. Going beyond these examples, Refs. [43, 44] remarkably show that for any constant $\alpha$, almost any function with growth $L^{2}\leq f(L)<L^{\alpha}$ is the Dehn function of some finitely presented semigroup; this includes for example “unreasonable-looking” functions such as $L^{7\pi},L^{2}\log L,$ and $L^{e}\log(L)^{\log(L)}\log\log L$. In addition to the worst case complexity of the word problem, we will also have occasion to consider its average-case complexity (both for groups and semigroups), a quantity has recieved much less attention in the math literature (the only exception the authors are aware of is [45]). To this end, we define the typical Dehn function $\mkern 1.5mu\overline{\mkern-1.5mu{\sf Dehn}\mkern-1.5mu}\mkern 1.5mu_{g}(L)$ as the number of steps needed to map a certain constant fraction of words in $K_{g,L}$ to one another:
@@ II.2.2 Space complexity
${\sf EL}(L)\triangleq{\sf EL}_{e}(L)\qquad\text{({\hbox{G}} a group)}$) (24)
as a simple metric of the space complexity. In App. B we show that ${\sf EL}(L)\lesssim L$ for all Abelian groups, and that all finite groups have ${\sf EL}(L)\leq L+C$ for some constant $C$; such scalings are “uninteresting” from the perspective of space complexity. Just as with the Dehn function though, there exist simple semigroups for which ${\sf EL}(L)$ grows extremely fast with $L$, the consequences of which will be explored in Sec. V.
@@ II.3 Semigroup dynamics and Hilbert space fragmentation
[…]
@@
For generic choices of $\lambda_{ab,i}$, the model is non-integrable, so any ergodicity breaking cannot be associated with integrability. Note that the product states $|w\rangle$ where $w=a_{1}\cdots a_{L}$ are annihilated by $H_{PF}$ if $a_{i}\neq a_{i+1}$ for all $i$. These product states alone provide $(2s+1)(2s)^{L-1}$ dynamically-disconnected dimension-1 sectors not attributable to any local symmetry, meaning that $H_{PF}$ exhibits HSF. The dynamics generated by $H_{PF}$ is in fact a special case of our construction applied to the group
@@
which matches the Hamiltonian in (26). Unfortunately the Dehn function of $\mathbb{Z}_{2}^{\ast(2s+1)}$ is easily seen to scale as ${\sf Dehn}(L)\sim L^{2}$, and thus does not provide a complexity scaling that is unusual enough to warrant further study. Having discussed the concept of HSF, let us further understand the structure and size of disconnected sectors under the dynamics ${\sf Dyn}_{G}$. The number of such disconnected sectors at least as large as the number of subspaces, each labelled by $K_{g,L}$ for some $g\in G$. Thus, the number of such subspaces—which we denote as $N_{K}(L)$—depends on the number of semigroup elements expressible as words of length $\leq L$. In particular, we define the geodesic length $|g|$ of an element $g\in G$ as the length of the shortest word representing $g$:
@@
A word $w$ with $\varphi(|w\rangle)=g$ satisfying $|w|=|g|$ is called a geodesic word. Then
@@
Semigroups with more elements “close” to the identity thus have dynamics with a larger number of Hilbert space fragments. As an example, one may readily verify that in the pair flip model, $N_{K}(L)\sim(2s)^{L}$ grows exponentially as long as $s>1/2$. For many models with group constraints, including the pair-flip model, $N_{K}(L)$ exactly determines the number of Hilbert space fragments. However, for some semigroups, $N_{K}(L)$ is not the full story, with each $K_{g,L}$ further fragmenting into subspaces in a way controlled by the expansion length function defined in (23). Understanding this phenomenon is the subject of Sec. V. In the study of HSF, a distinction is often made between “strong” and “weak” HSF. This distinction was originally discussed in the context of Hamiltonian dynamics [21], where it was defined by violations of strong and weak ETH, respectively. Since we are discussing things at a level where the nature of ${\sf Dyn}_{G}$ may or may not involve eigenstates, we will instead adopt a slightly different definition in terms of the size of the largest $K_{g,L}$ sector, which we denote as $K_{{\rm max},L}$ (in App. B we prove that for groups, the largest sector is in fact always the one associated with the identity, $K_{{\rm max},L}=K_{e,L}$). We will say that the dynamics is
@@
We will find it useful to subdivide the strongly fragmented case into additional classes according to how quickly $|K_{{\rm max},L}|/|\mathcal{H}|$ vanishes as $L\rightarrow\infty$. We will say that the fragmentation in the case of strong HSF is
@@
Note that the above definitions are made without reference to any global symmetry sectors. This means that if global symmetries happen to be present, the quantum numbers associated with them will constitute part of the elements $g$ labeling the different dynamical sectors, and the above definition of strong/weak HSF will simply single out the largest, regardless of its symmetry quantum number(s). When symmetries are present one could also quantify the degree of fragmentation by first fixing a quantum number, changing $K_{\rm max}(L)$ to be the largest sector having that quantum number, and replacing $|\mathcal{H}|$ by the dimension of the chosen symmetry sector. However, since generic ${\sf Dyn}_{G}$ dynamics needn’t have any global symmetries, and since the result of the above procedure can depend sensitively on the chosen quantum number [48, 49, 50], we will focus only on the above (simpler) definition, which maximizes over all symmetry sectors. Our models provide a way of addressing two questions raised in Ref. [21]. The first question was whether or not 1D models exist with $10<|K_{{\rm max},L}|/|\mathcal{H}|<$ 1 in the thermodynamic limit (known examples with weak fragmentation all have $|K_{{\rm max},L}|/|\mathcal{H}|\rightarrow 1$ in the thermodynamic limit). Our construction answers this question in the affirmative, with examples provided by ${\sf Dyn}_{G}$ for any finite non-Abelian $G$ (e.g. $G=S_{3}$). The second question concerned the existence of models where $|K_{{\rm max},L}|/|\mathcal{H}|$ vanishes more slowly than exponentially as $L\rightarrow\infty$ after specifying to a fixed symmetry sector (in fact an affirmative answer to this question was already provided by the spin-1 Motzkin chain introduced in [51], where $|K_{{\rm max},L}|/|\mathcal{H}|\sim L^{-3/2}$ [^11] This follows from the fact that the probability of a length-$L$ random walk on $\mathbb{Z}$ being a returning Motzkin walk scales as $\sim L^{-3/2}$.). Our models provide (many) more examples of this phenomenon, as one need only let $G$ be a group with $L^{2}\lesssim{\sf Dehn}(L)\lesssim L^{\infty}$, a simple example being the discrete Heisenberg group $\mathsf{H}_{3}$. [^12] If ${\sf Dyn}_{G}$ possesses global symmetries and one modifies the above definitions to operate within a fixed symmetry sector, then various models are known where a specific symmetry sector with polynomially strong fragmentation exists [48, 49, 50]. One may further ask whether there are strongly fragmented systems where $|K_{{\rm max},L}|/|\mathcal{H}|$ decays at a rate in between $1/{\rm poly}(L)$ and $\exp(-L)$. The answer to this is again affirmative, with one such example being the focus of Sec. IV.